OpenStudy (anonymous):

what is y=5*5/n metal and fan

OpenStudy (anonymous):

y= 25 5+5=10 10+5=15 15+5=20 20+5=25

OpenStudy (anonymous):

did you divde

OpenStudy (anonymous):

No, i did 5*5 to check you answer you can do 25/5

OpenStudy (anonymous):

but y=25

OpenStudy (anonymous):

theres your medal

OpenStudy (anonymous):

thank you